OpenStudy (anonymous):

3.6 million people voted in the last state-wide election for governor. Governor Smith won the election with 52.1% of the votes. About how many people voted for Governor Smith?

OpenStudy (ranga):

3.6 x 10^6 x 52.1 / 100 = ?

OpenStudy (anonymous):

What is 52.1% of 3600000 = 52.1% * 3600000 = 0.521 * 3600000 = 1875600

OpenStudy (anonymous):

52.1% = 52.1 / 100 = 0.521

OpenStudy (ranga):

A million is 10^6

OpenStudy (anonymous):

1875600 people voted for smith

OpenStudy (ranga):

\[\frac{ 3.6 \times 10^6 \times 52.1 }{ 100 } = 1,875,600\]